Session: Nov 2001 (session 3)
Wavelength: 5.0 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (orfei at Tue Nov 6 12:13:01 MET 2001) Success

Jodrell_Bank (Alastair Gunn at Tue Nov 6 12:28:47 MET 2001) Success
Phase cal signal too strong so only switched on intermittently. No other known problems.

Cambridge (Alastair Gunn at Tue Nov 6 12:31:28 MET 2001) Success
Signal levels in IF reduced due to upconvertor level too low. Levels increased towards end of experiment. No other known problems.

Effelsberg (A. Kraus at Tue Nov 6 13:32:33 MET 2001) Failure
Effelsberg could not participate.

Torun (K.Borkowski at Wed Nov 7 10:46:27 MET 2001) Failure
Could not write to tape with the Penny&Giles recorder upgraded to two heads and FS 9.5 (an FS/recorder problem).

Onsala (John Conway at Mon Nov 12 09:47:57 MET 2001) Minor failures
Observed despite high winds near limit for telescope pointing will not be good.